第二部分:概率与统计 [60分]

  1. (a)

    若 P(B∣A)=0.2\text{P}(B \mid A) = 0.2,P(A∣B)=0.6\text{P}(A \mid B) = 0.6 且 P(A′∩B′)=0.3\text{P}(A' \cap B') = 0.3,求 P(A∩B)\text{P}(A \cap B)。

  2. (b)

    事件 XX、YY 和 ZZ 满足事件 XX 和 ZZ 互斥。已知 P(X∩Y)=0.1\text{P}(X \cap Y) = 0.1,P(X′∩Y)=0.35\text{P}(X' \cap Y) = 0.35,P(Y∩Z)=0.2\text{P}(Y \cap Z) = 0.2 且 P(X∪Y∪Z)=0.95\text{P}(X \cup Y \cup Z) = 0.95,求 P(X)\text{P}(X) 的最小值和最大值。

先不看答案,让 AI 诊断你卡在哪里?

它不会直接告诉你答案,而是问你卡在哪一步看看 AI 是怎么陪你一步步弄懂这道题的

完整解法

6(a)

1. Express $\text{P}(A)$ and $\text{P}(B)$ in terms of $\text{P}(A \cap B)$ using the conditional probability formulas.

P(B∣A)=P(A∩B)P(A)  ⟹  P(A)=P(A∩B)0.2=5P(A∩B)\text{P}(B \mid A) = \frac{\text{P}(A \cap B)}{\text{P}(A)} \implies \text{P}(A) = \frac{\text{P}(A \cap B)}{0.2} = 5\text{P}(A \cap B)
P(A∣B)=P(A∩B)P(B)  ⟹  P(B)=P(A∩B)0.6=53P(A∩B)\text{P}(A \mid B) = \frac{\text{P}(A \cap B)}{\text{P}(B)} \implies \text{P}(B) = \frac{\text{P}(A \cap B)}{0.6} = \frac{5}{3}\text{P}(A \cap B)

2. Use the complement rule and De Morgan's laws to evaluate $\text{P}(A \cup B)$.

P(A′∩B′)=1−P(A∪B)\text{P}(A' \cap B') = 1 - \text{P}(A \cup B)
0.3=1−P(A∪B)  ⟹  P(A∪B)=0.70.3 = 1 - \text{P}(A \cup B) \implies \text{P}(A \cup B) = 0.7

3. Use the addition rule to form an equation and solve for $\text{P}(A \cap B)$.

P(A∪B)=P(A)+P(B)−P(A∩B)\text{P}(A \cup B) = \text{P}(A) + \text{P}(B) - \text{P}(A \cap B)
0.7=5P(A∩B)+53P(A∩B)−P(A∩B)0.7 = 5\text{P}(A \cap B) + \frac{5}{3}\text{P}(A \cap B) - \text{P}(A \cap B)
0.7=(4+53)P(A∩B)0.7 = \left( 4 + \frac{5}{3} \right) \text{P}(A \cap B)
0.7=173P(A∩B)0.7 = \frac{17}{3} \text{P}(A \cap B)
P(A∩B)=0.7×317=2.117=21170\text{P}(A \cap B) = \frac{0.7 \times 3}{17} = \frac{2.1}{17} = \frac{21}{170}

答案

21170\frac{21}{170}

6(b)

1. Find $\text{P}(Y)$ using the given intersections.

The event YY can be partitioned into two mutually exclusive parts, X∩YX \cap Y and X′∩YX' \cap Y:
P(Y)=P(X∩Y)+P(X′∩Y)\text{P}(Y) = \text{P}(X \cap Y) + \text{P}(X' \cap Y)
P(Y)=0.1+0.35=0.45\text{P}(Y) = 0.1 + 0.35 = 0.45

2. Express $\text{P}(X \cup Y \cup Z)$ using the addition rule to form an equation linking $\text{P}(X)$ and $\text{P}(Z)$.

Using the addition rule on the events (X∪Z)(X \cup Z) and YY:
P(X∪Y∪Z)=P(X∪Z)+P(Y)−P((X∪Z)∩Y)\text{P}(X \cup Y \cup Z) = \text{P}(X \cup Z) + \text{P}(Y) - \text{P}((X \cup Z) \cap Y)
Since XX and ZZ are mutually exclusive (P(X∩Z)=0\text{P}(X \cap Z) = 0), we have P(X∪Z)=P(X)+P(Z)\text{P}(X \cup Z) = \text{P}(X) + \text{P}(Z).
Furthermore, the intersection with YY gives (X∪Z)∩Y=(X∩Y)∪(Z∩Y)(X \cup Z) \cap Y = (X \cap Y) \cup (Z \cap Y). Because X∩Z=∅X \cap Z = \emptyset, these two components are also mutually exclusive:
P((X∩Y)∪(Z∩Y))=P(X∩Y)+P(Y∩Z)\text{P}((X \cap Y) \cup (Z \cap Y)) = \text{P}(X \cap Y) + \text{P}(Y \cap Z)
=0.1+0.2=0.3= 0.1 + 0.2 = 0.3
Substitute these into the addition rule equation:
0.95=(P(X)+P(Z))+0.45−0.30.95 = (\text{P}(X) + \text{P}(Z)) + 0.45 - 0.3
0.95=P(X)+P(Z)+0.150.95 = \text{P}(X) + \text{P}(Z) + 0.15
P(X)+P(Z)=0.8\text{P}(X) + \text{P}(Z) = 0.8

3. Establish bounds for $\text{P}(X)$ based on the minimum possible probabilities of subsets.

Since a subset's probability cannot exceed the whole event's probability (and probabilities must be non-negative):
X∩Y⊆X  ⟹  P(X)≥P(X∩Y)=0.1X \cap Y \subseteq X \implies \text{P}(X) \ge \text{P}(X \cap Y) = 0.1
Y∩Z⊆Z  ⟹  P(Z)≥P(Y∩Z)=0.2Y \cap Z \subseteq Z \implies \text{P}(Z) \ge \text{P}(Y \cap Z) = 0.2

Substitute P(Z)=0.8−P(X)\text{P}(Z) = 0.8 - \text{P}(X) into the second inequality:
0.8−P(X)≥0.20.8 - \text{P}(X) \ge 0.2
P(X)≤0.6\text{P}(X) \le 0.6

Combining the bounds, we have:
0.1≤P(X)≤0.60.1 \le \text{P}(X) \le 0.6

答案

Minimum value =0.1= 0.1, Maximum value =0.6= 0.6

AI 分析,可能出错

用 AI 批改,圈出写错的那一步,并自动将你的薄弱点存入错题本,形成你的错因库。

同一知识点的其他题

  1. 1已知事件 AA 和 BB 满足 P(B)=13\mathrm{P}(B) = \frac{1}{3},P(A∣B)=25\mathrm{P}(A|B) = \frac{2}{5} 且 P(A∪B)=35\mathrm{P}(A \cup B) = \frac{3}{5}。…
  2. 2凯(Kai)和里奥(Leo)两兄弟是定期训练的双项全能(跑步和骑行)运动员。在每一天中,凯骑行的概率为 34\frac{3}{4},跑步的概率为 35\frac{3}{5},两者都进行的概率为 pp。 平均而言,里奥在一周的 7 天中有…
  3. 3一对同胞 Abbey 和 Betty 还有其他六个朋友一起去参加书展。他们八个人随机排成一队等候一位著名作家的签名。求满足以下条件的概率: Abbey 排在队列第一位且 Betty 排在队列第六位, Abbey 排在队列第一位或 Betty…
  4. 4已知某种特定疾病的感染率约为万分之一(每 10 000 人中约有 1 人感染)。针对该疾病的一项检测准确率为 99.9%,即检测结果与真实情况相符的概率为 99.9%。求在已知某人检测结果呈阳性的条件下,此人确实患有该疾病的概率。 研究人员…
  5. 5对于事件 AA 和 BB,已知 P(A)=0.55\mathrm{P}(A) = 0.55,P(B)=0.65\mathrm{P}(B) = 0.65 且 P(A∪B)=0.93\mathrm{P}(A \cup B) = 0.93。 求 P(A∩B)\mathrm{P}(A \cap B)…